Pronghorn Company offers an MP3 download (seven-single medley) as a premium for every 6 candy bar wrappers presented by customers together with $3.40. The candy bars are sold by the company to distributors for 30 cents each. The purchase price of each download code to the company is $3.15. In addition, it costs 50 cents to distribute each code. The results of the premium plan for the years 2017 and 2018 are as follows. (All purchases and sales are for cash.)

2017

2018

MP3 codes purchased 345,000 455,400
Candy bars sold 2,998,000 2,809,300
Wrappers redeemed 1,656,000 2,070,000
2017 wrappers expected to be redeemed in 2018 400,200
2018 wrappers expected to be redeemed in 2019

483,000

a) Prepare the journal entries that should be made in 2017 and 2018 to record the transactions related to the premium plan of the Pronghorn Company.

b) Indicate the amounts for each accounts, and classifications of the items related to the premium plan that would appear on the balance sheet and the income statement at the end of 2017 and 2018.

Amount

Account

2017

2018

Classification

Inventory of Premiums $
Premium Liability
Premium Expense
Answer #1

a)

No. General Journal Debit Credit
2017
1 Inventory of Premiums (345000 x $3.15) 1086750
Cash 1086750
(To record the purchase of premium inventory)
2 Cash 899400
Sales Revenue (2998000 x $0.30) 899400
(To record the sales)
3 Cash [1656000/6 = 276000 x ($3.40 - $0.50)] 800400
Premium Expense [276000 x ($3.15 + $0.50 - $3.40)] 69000
Inventory of Premiums (276000 x $3.15) 869400
(To record the wrappers redeemed)
4 Premium Expense 16675
Premium Liability [400200/6 = 66700 x ($3.15 + $0.50 - $3.40)] 16675
(To record the premium liability)
2018
1 Inventory of Premiums (455400 x $3.15) 1434510
Cash 1434510
(To record the purchase of premium inventory)
2 Cash 842790
Sales Revenue (2809300 x $0.30) 842790
(To record the sales)
3 Cash [2070000/6 = 345000 x ($3.40 - $0.50)] 1000500
Premium Liability 16675
Premium Expense [(345000 - 66700) x ($3.15 + $0.50 - $3.40)] 69575
Inventory of Premiums (345000 x $3.15) 1086750
(To record the wrappers redeemed)
4 Premium Expense 20125
Premium Liability [483000/6 = 80500 x ($3.15 + $0.50 - $3.40)] 20125
(To record the premium liability)

b)

Account Classification Amount
2017 2018
Inventory of Premiums Balance Sheet - Current Asset 217350 565110
Premium Liability Balance Sheet - Current Liability 16675 20125
Premium Expense Income Statement - Selling Expenses 85675 89700